# Active Inference AI Systems for Scientific Discovery

### v3

@@ 5   Summary and Outlook
This perspective calls upon the community to build on substantial progress in causal machine learning, active learning, and automated scientific discovery to address these critical gaps. The causal machine learning community has made significant strides in developing methods for causal inference from observational data, with frameworks such as Pearl’s causal hierarchy and recent advances in causal representation learning providing mathematical foundations for understanding interventions and counterfactuals. Similarly, active learning has evolved sophisticated strategies for optimal experimental design, while automated discovery systems have demonstrated success in specific domains such as materials science and drug discovery. However, these communities have largely operated in isolation, with causal methods focusing primarily on statistical inference rather than physical mechanism discovery, active learning optimizing for narrow uncertainty reduction rather than conceptual breakthroughs, and automated discovery systems excelling at interpolation within known spaces rather than extrapolation to genuinely novel phenomena. Realizing such a capability requires a diverse consortium to co-develop models, agentic infrastructure, formal verification stacks, simulators, robotic labs, and evaluations. Success cannot merely be measured by benchmarks alone, but by the moment when these systems become truly useful –as judged by domain experts– and make genuinely novel scientific discoveries.
@@ Acknowledgment
A comprehensive review of agentic systems for scientific discovery can be found in Ref. [24]. Below, a few references that are related to abstraction, reasoning and reality gaps are provided.
@@
These implementations and others are already accelerating science, but also collectively reveal a critical insight: current systems excel at automating well-defined scientific workflows but falter when required to navigate the uncertain terrain of genuine discovery. They can execute sophisticated experimental protocols, analyze complex datasets, and even generate plausible hypotheses, yet they lack the metacognitive capabilities to recognize when they are operating beyond their training domains.

### v4

@@ 5   Outlook
Active AI systems encompass external experience (empirical data) and internal experience (mental simulation). AI systems that can fluidly navigate between these modes will mark the transition from tools that find patterns to partners that discover principles. This perspective builds upon substantial progress in causal machine learning, active learning, and automated scientific discovery while addressing critical gaps. The causal machine learning community has made significant strides in developing methods for causal inference from observational data, with frameworks like Pearl’s causal hierarchy and recent advances in causal representation learning providing mathematical foundations for understanding interventions and counterfactuals [86, 52]. Similarly, active learning has evolved sophisticated strategies for optimal experimental design [88, 35], while automated discovery systems have demonstrated success in specific domains such as materials science and drug discovery [15, 91]. However, these communities have largely operated in isolation, with causal methods focusing primarily on statistical inference rather than physical mechanism discovery, active learning optimizing for narrow uncertainty reduction rather than conceptual breakthrough, and automated discovery systems excelling at interpolation within known spaces rather than extrapolation to genuinely novel phenomena. We note that efforts to bridge these communities are underway: the Acceleration Consortium [95] represents a major initiative integrating autonomous laboratories with AI-driven discovery, while recent work demonstrates promising integration of causal reasoning with materials optimization [44, 31, 16].
Current implementations prioritize task completion over understanding, optimization over exploration, and correlation over causation. The path forward requires AI systems that integrate causal reasoning not merely as a statistical tool but as the foundation for mental simulation and counterfactual experimentation, extending active learning beyond data efficiency to include the generation and testing of novel hypotheses that violate existing assumptions, and grounding automated discovery in continuous empirical feedback loops that prevent drift from physical reality. Most critically, while existing approaches excel within their prescribed domains, they lack the architectural foundation for the kind of open-ended, cross-domain reasoning that characterizes human scientific discovery—the ability to recognize when anomalous observations demand not just parameter updates but fundamental reconceptualization of the problem space itself.
Scientific discovery has always been a collaborative enterprise—across disciplines, institutions, and generations. AI systems represent new kinds of collaborative tools. The transition from static models to living systems marks a fundamental shift in how we conceive of AI systems that persist, that remember, that build intuition through repeated engagement with reality. Just as human scientists develop insight through years of experimentation, future AI systems will accumulate wisdom through continuous cycles of hypothesis, experiment, and revision. We thus call for the creation of new benchmarks and research programs centered around the proposed stacked architecture, moving evaluation beyond static datasets to interactive, discovery-oriented environments. Success should ultimately be judged not only by benchmark performance but by domain expert assessment of discovery quality. For example, whether AI-generated hypotheses lead to experiments that domain scientists find surprising and informative, whether the system identifies gaps in current understanding that experts recognize as important, and whether it proposes experimental designs that are both feasible and scientifically meaningful. Structured evaluation protocols, such as blinded assessment of AI-proposed versus human-proposed hypotheses by independent expert panels, could provide more rigorous qualitative evaluation [69].
Finally, it has to be emphasized that modern AI systems are already useful in their present form, and are being utilized effectively by scientific research groups across the world. However, even with future improvements, these tools bring many systemic hazards [66]: a) false positives and false negatives: spurious correlations can be mistaken for laws, while cautious priors may hide real effects, and thus rigorous uncertainty metrics and adversarial falsification must be built in; b) epistemic overconfidence: large models can exhibit poorly calibrated uncertainty estimates, particularly on inputs far from the training distribution, a phenomenon where predictive confidence remains high even when accuracy degrades [40]; ensemble disagreement among multiple independently trained models provides one diagnostic for such failures; c) erosion of insight and rigor: over time, there is signficiant risk of researchers losing key scientific skills; d) Cost: simulation-driven exploration can consume resources long after the expected reduction in uncertainty from additional experiments becomes negligible relative to their cost (i.e., marginal information gain saturates); resource allocation algorithms (schedulers) must weigh value against resources; e) instrumental drift: scientific instruments and sensors evolve over time, and their characteristics may shift due to wear, recalibration, or environmental changes; this can be addressed through standard calibration practices such as round-robin testing across laboratories [108], but without continual residual checks and rapid retraining, predictions may silently bias. These issues have to be continually acknowledged, recognizing and safeguards should be embeded into the scientific process.
@@ Acknowledgment
A comprehensive review of agentic systems for scientific discovery can be found in Ref. [38]; see also Zenil et al. [106] for a complementary perspective. Below, a few references that are related to abstraction, reasoning and reality gaps are provided.
Before the transformer era, several systems demonstrated that AI could make genuine scientific discoveries through closed-loop interaction with physical laboratories. Adam and Eve [54, 100] combined ontological reasoning with active learning to autonomously design and execute falsifiable experiments. More recently, the Robot Chemist [14] demonstrated autonomous exploration of reaction conditions, while A-Lab [92] achieved autonomous synthesis of novel materials. These systems embody key principles of our proposed architecture: persistent scientific memory (through ontologies and databases), closed-loop empirical validation, and hypothesis-driven experimentation. However, they operate within well-defined search spaces with predetermined objectives, limiting their capacity for the kind of open-ended discovery that might restructure the problem space itself.
@@
These implementations and others are already accelerating science, but also collectively reveal a critical insight: current systems excel at automating well-defined scientific workflows but falter when required to navigate the uncertain terrain of genuine discovery. They can execute sophisticated experimental protocols, analyze complex datasets, and even generate plausible hypotheses, yet they lack the metacognitive capabilities to recognize when they are operating beyond their training domains. The contribution of this perspective is to identify the architectural requirements for moving beyond these limitations, not to claim that closed-loop discovery is impossible, but to argue that scaling to genuinely novel discovery requires integrated resolution of the abstraction, reasoning, and reality gaps.